OpenStudy (anonymous):

The number of messages sent to a computer bulletin board is a Poisson distributed random variable with a mean of 5 messages per hour. Round your answers to four decimal places (e.g. 98.7654). (a) What is the probability that 5 messages are received in 1 hour? (b) What is the probability that 10 messages are received in 1.5 hours? (c) What is the probability that less than two messages are received in one-half hour? Can anyone help me with this ?

OpenStudy (anonymous):

@wio hey can u help me with this question ?

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

\[ \Pr(X=k) = \frac{\lambda^ke^{-\lambda }}{k!} \]

OpenStudy (anonymous):

The mean is \(\lambda=5\)

OpenStudy (anonymous):

(1) asks for \(\Pr(X = 5)\)

OpenStudy (anonymous):

For (2) you as talking about 1.5 hours, so I believe you need to scale lambda.\[ \lambda' = 1.5\lambda = 1.5(5)=7.5 \]

OpenStudy (anonymous):

And then use \(k=10\).

OpenStudy (anonymous):

For (3) we have \[ \lambda ''=\frac 12 \lambda =\frac 12 (5) = 2.5 \]And you want to find probability for \( k=0,1 \) individually and add them up

OpenStudy (anonymous):

Thank you so much for your help! it was very clear and helpful :)